# The double gamma function and Vladimir Alekseevsky

### v2

@@
Considering the Weierstrass product for the entire function $\sin\pi x$ and taking half of the factors corresponding to non-positive roots, we obtain the function $1/\Gamma(x)$. Considering the Weierstrass product for the Jacobi theta function $\vartheta_{1}$ and taking quarter of the factors (corresponding to the positive quadrant in the lattice $n_{1}\omega+n_{2}\omega_{2}$ of periods), we come to the double gamma function. It satisfies big collection of identities parallel to properties of $\Gamma(x)$. We discuss the seminal work (1888-89) by V. P. Alekseevsky, where the double gamma function was introduced and investigated. His standpoint was the paper by P. É. Appell (1881) on extensions of Heine’s $q$-version of the Euler gamma function. For the case of equal periods $\omega_{1}=\omega_{2}$, a function equivalent to the double gamma function was introduced earlier by H. Kinkelin (1860). After Alekseevsky, the double gamma function was a topic of investigations by E. W. Barnes, J. Beaupin, G. H. Hardy, V. A. Steklov in 1899-1907. It was quite unusual compared to special functions known at that time. Despite obviously bright publications, the function was forgotten until 1976-1979, when it appeared again in works by T. Shintani and M. V. Vignéras. In the same time, Shintani introduced the double sine as a product of two double gammas. We also discuss the biography of Alekseevsky (1858-1916).
@@
Key words. Mathematics of XIX century, special functions, gamma function, theta function, complex analysis, functional equations, Weierstrass factorization theorem, universities, biography.
1. The work by Alekseevsky, 1888
@@
Alekseevsky’s [^1] Vladimir Petrovich Alekseevsky (1858-1916), Russian: Алексеевский; old Russian orthography: Алексевскй. We use the most natural English spelling of the name, German and French transliterations used in 1884-1907 were Alexeiewsky, Alexejewsky, Alexéievsky. work On functions similar to the gamma function was published in Communications of Kharkov mathematical society [^2] It was a well-known journal during 1879-1918. Russian: Сообщения и протоколы заседаний Математического общества при Императорском Харьковском университете; old Russian orthography: Сообщеня и протоколы засданй Математическаго общества при Императорскомъ Харьковскомъ университет; French: Communications et procus-verbaux de la société mathématique de Kharkow; German: Sammlung der Mitteilungen und Protokolle der Mathematischen Gesellschaft in Charkow. in two papers [Ale1888], [Ale1889], 1888-1889.
@@
1.1. The function $G(x)$. The article by Alekseevsky begins with the words: The first problem, whose solution is the topic of the present investigation, is to investigate properties of a function $G(x)$ satisfying the equation:
@@
Here there are no references, no names, no motivations, no acknowledgements. It is hard to believe that the author simply decided to attack such an exotic problem and successfully solved it. Actually, the exposition in the article was backwards to the real way of thinking; this is discussed below in Subsect. 1.
We pass to logarithms in (1.1) and come to the difference equation
@@
According Malmsten formula, see, e.g., [Erd1953], formula 1.9(1),
@@
Denote by $L_{1}(x,u)$ the integrand. Alekseevsky (§1) writes the simplest function $L_{2}(x,u)$ such that
@@
is the Euler constant (cf. the similar expression for $\Gamma(x)$ in [Erd1953], 1.1 (3)). Next, he noticed that this formula gives an expression for $G(x)$ on the whole complex plane.
@@
After complicated calculations with functions $G(x;\alpha)$ in §21 (see below Subsect. 1) he gets the multiplication formula (cf. [Erd1953], 1.2(11)) in the form
@@
The multipication formula admits a simpler form obtained by Kinkelin [Kin1860], see below formula (2.2).
In §11, Alekseevsky establishes the following counterpart of ’Stirling’ asymptotic expansion (cf. [Erd1953], 1.18(1)):
@@
In §16, there are introduced functions $G_{n}(x)$ satisfying the relations
@@
and this implies the desired property.
Alekseevsky does not develop further theory of such functions, but in §17-19 discusses some generalities of function theory related to the Weierstrass factorization theorem. Namely, he considers arbitrary products of the form
@@
We represent $\ln f_{0}(z)$ as a Laplace transform, and, applying the same trick with functions $P_{n}(z)$, construct sequence of functions $f_{1}(z)$, $f_{2}(z)$, …such that
$f_{n}(z+1)=f_{n}(z)f_{n+1}(z)$
Alekseevsky write explicit representations of functions $f_{n}$ as products of functions $G_{m}$ with given $m\leqslant n$.
1.4. The double gamma functions with two periods. In §20, Alekseevsky writes: Functions similar to $\Gamma(x)$ considered above are special cases of more general functions. Since they are related to theta-functions, we derive some properties of one of such functions.
Fix a parameter $\alpha$ (which initially is positive real, but later it becomes complex). Alek-seevsky considers a function [^3] He uses the notation $H(x)$ or $H(x,\alpha)$. $G(x;\alpha)$ such that
@@
Alekseevsky finds main properties (cf. [AlKu2023]) of this function, in particular,
@@
see §21. This formula has numerous interesting special cases, for instance, we can set $m=1$, or $n=1$, or $\alpha=1$, or $\alpha=1/n$, or $\alpha=n$, or $m=n$. Multiplication formula (1.8) also is a non-obvious corollary of the last formula.
@@
It is clear that this function consists of quarter factors entering to $\varsigma_{3}$-function of Weierstrass [^4] The Weierstarss function, see, e.g., [Akh1970], §13, is given by a product (1.19) over $(m,n)\in\mathbb{Z}^{2}\setminus(0,0)$. In our case (1.16), the product is given over $m\geqslant 0$, $n\geqslant 0$, $(m,n)\neq(0,0$. (or $\vartheta_{1}$ of Jacobi), herein it comes from $\varsigma_{3}$ in the same way as $\frac{1}{\Gamma(z)}$ from $\sin\pi z$. However, in a two-periodic function, a ratio $\frac{\omega^{\prime}}{\omega}$ can be imaginary or complex but for the function $G$ this condition is not necessary555i.e., a ratio of periods can be positive real..
The function $G$ is analogous to the function of Heine [^6] Footnote from [Ale1888]: Handbuch der Kugelfunctionen, p. 109, or Crelle J. t 34. p. 290, see [Hei1847], [Hei1878]. I do not understand these references (they are taken from [App1881]). However, the attribution of this expression to Eduard Heine is correct, see the next subsection., which is composed from half of the factors of $\vartheta_{3}$. For a collation, we present an expression from Appell’s memoire [^7] Footnote from Alekseevsky: Mathematische Annalen. B. 19. p. 84. , see [App1881],
@@
This explains the origin of this work. Looking to the formula in Appell’s paper, Alekseevsky understood that it is possible to manipulate with Weierstrass-type product over a quadrant in the lattice $\mathbb{Z}^{2}$. However, he found that the integral representation (1.4) is a key to the whole story and the initial definition of the work became the final theorem of the published paper (in fact, an evaluation of constants $a$, $b$ in (1.16) is highly nontrivial question).
The work under discussion became a dissertation of Alekseevsky [Ale1892]. It was published as separate book and is a literal reprint of [Ale1888]-[Ale1889] with an addition of 4 pages in the beginning. The introduction contains the following sentence:
The known dependence between the Euler function of the second time [the gamma function] and the sine [i.e., the reflection formula $\Gamma(z)\Gamma(1-z)=\pi/\sin(\pi z)$] made me think to extract from the function $\vartheta_{1}$ of Jacobi a function similar to $1/\Gamma(x)$ and investigate its properties.
By way of analogy with the function $\Gamma(x)$, I came to the consideration of functions whose roots have the form $m\omega+n\omega^{\prime}$, where $m$ and $n$ are negative integers, and $\omega$, $\omega^{\prime}$ are periods of elliptic functions.
As we have seen above, Alekseevsky introduced additional parameters $\omega_{1}$, $\omega_{2}$ to the function $G=G_{2}(x)$. After Barnes [Bar1904], we know that it is possible to introduce similar parameters $\omega_{1}$, …, $\omega_{n}$ to the functions $G_{n}(x)$ and get ‘multigamma functions, such functions (we use a non-standard notation) satisfy the collection of identities
@@
It is strange that Appell himself did not introduce the double gamma. In any case, the response of Alekseevsky to the Appell’s paper was completely natural.
@@
2. Roots of functions similar to $\Gamma$ has a form

### v3

@@
This paper is about a forgotten function and a forgotten mathematician. The double gamma function is now an important special function, which appears for different reasons in many branches of mathematics and in mathematical physics, as it satisfies a large and amusing collection of identities parallel to the classical gamma function. It was discovered and investigated in detail by Vladimir Alekseevsky in 1888-89. We outline his starting point here: considering the Weierstrass product for the entire periodic function $\sin\pi x$ and taking half of the factors corresponding to non-positive roots of the sine, we obtain the function $1/\Gamma(x)$. Considering the Weierstrass product for the doubly quasiperiodic Jacobi theta function $\vartheta_{1}$ and taking the half of factors we come to the $q$-gamma function (which appears, with a slight change in notation, in Eduard Heine’s work). Considering the quarter of the factors (corresponding to the positive quadrant in the lattice $n_{1}\omega+n_{2}\omega_{2}$ of periods), Alekseevsky arrived at his double gamma function. The work in this direction was continued by Ernest Barnes, Jean Beaupin, Godfrey Hardy, and Vladimir Steklov in 1899–1907. After 1907, publications on this topic stopped, and the double gamma appeared again 70 years later due to Takuro Shintani. In this paper we discuss Alekseevsky’s seminal work and its genesis, the history of the double gamma, and Alekseevsky’s biography (1858-1916).
@@
Key words. Mathematics of XIX century, mathematics of XX century, special functions, gamma function, theta function, complex analysis, functional equations, Weierstrass factorization theorem, universities, biography.
4000000000000000000000000.Content.
1. Introduction. A forgotten function and a forgotten mathematician
2. Biography of Alekseevsky
3. Alekseevsky’s work, 1888-89
4. On the history of the double gamma function
5. Alekseevsky: historiography, lost work, and publications
6. Conslusion.
4.References
1. Introduction. A forgotten function and a forgotten mathematician
In 1996, Vladimir Fock spoke to me about the work [Fad1995], [Fock1997], [Kash1998], [FoCh1999] being done on the quantization of Teichmüller spaces; this construction was essentially based on a couple amusing special functions called ’Barnes double sine’ and ’Barnes double gamma’. I read the original papers of Ernest Barnes [Bar1899][Bar1904] and observed that he referred to an earlier work of Alexeiewsky. This name was unknown to me, and my immediate attempt to find an information about this person and his works was unsuccessful.
Many years later, collecting material for a book about Nikolay Luzin, I opened Krulikovsky’s ’History of the development of mathematics in Tomsk[Kru1967]. Here I was again met with the name Aleksevsky. I found his main work and was very surprised reading it. I then started a historical excavation, which I now present to the reader.
The double gamma function is now a well recognized special function used in different branches of mathematics and in mathematical physics. It is an entire function $G(x;\alpha)$ of a complex variable $x$ and depending on a complex parameter $\alpha$, which satisfies the following functional relations:
(1.1) $G(x+1;\alpha)$ $=\Gamma\Bigl(\frac{x}{\alpha}\Bigr)\,G(x;\alpha);$
(1.2) $G(x+\alpha;\alpha)$ $=(2\pi)^{\frac{\alpha-1}{2}}\alpha^{-\frac{2x-1}{2}}\,\Gamma(x)\,G(x;\alpha),$
and is normalized by
$G(0;\alpha)=1.$
This function satisfies a big collection of properties parallel to the classical gamma function.
A special case of $G(x,\alpha)$ corresponding to $\alpha=1$ was discovered by Hermann Kinkelin in [Kin1860]; notice that in this case the two equations (1.1), (1.2) coincide. His remarkable work did not attract interest and was forgotten for a long time. The double gamma for general $\alpha$ was introduced by Vladimir Alekseevsky [^1] Russian: Алексеевский; old Russian orthography: Алексевскiй. I use the most natural English spelling of the name, which is used by his modern homonyms. German and French transliterations used in 1884-1907 were Alexeiewsky, Alexejewsky, Alexéievsky. in a large memoir [Ale1888][Ale1889]. This work was continued by Ernest Barnes [Bar1899], [Bar1900.1], [Bar1900.2], [Bar1901], [Bar1903], [Bar1904], Jean Beaupain [Bea1903], [Bea1904], [Bea1907], Godfrey Hardy [Har1905.1], [Har1905.2], and Vladimir Steklov [Ste1904]. The chain of publications on the double gamma and the multiple gamma was interrupted in 1907, and after which they were more or less forgotten for 70 years.
Alekseevsky (1858–1916), the discover of the double gamma, was also forgotten. He was mentioned in a few Russian papers on the histories of mathematics and education, [Gus1950], [Gus1952], [Kru1967], [Bakh1967], [Kru1967.1], [Yush1968], [Gran2003], [Bel2000] (mainly, on the local history of mathematics in Kharkov and Tomsk). He studied and worked at Kharkov University (1877-1905) and Tomsk Technological Institute (1905-1916). For the Kharkov period, we mainly have records of unsuccessful attempts to advance his career. In Tomsk he became a full professor of mechanics, and in 1907-1911 he was the director (rector) of the Institute.
We briefly discuss his biography in Sect. 2. But our main topics are his impressive central work and the difficulty of mathematical development.
In Sect. 3, we discuss Alekseevsky’s main work [Ale1888][Ale1889] and its genesis. It was a non-obvious response to Eduard Heine’s investigations of $q$-hypergeometric functions (though, more precisely, Alekseevsky’s starting point was Paul Appell’s work [App1881] on extending the $q$-gamma).
In Sect. 4, we discuss the history of the double gamma. We start with a brief exposition of Kinkelin’s work. It was basically surpassed in [Ale1888][Ale1889], but it also contains important results, which were not repeated in Alekseevsky’s memoir (the link between Kinkelin and Alekseevsky was later observed by Beaupin). We also try to understand the origin of Kinkelin’s work. It seems to have been an investigation in a ’speleological’ style, which is relatively common in theory of special functions.
The forgetting of important research [^2] Opening the database MathSciNet, which is quite selective, I see 233 references to [Bar1899] starting 2002. At the very least, this is unusual for a paper published a hundred years before. Google Scholar shows 512 references to [Bar1904]. is in itself an interesting historical phenomenon. In this case, we see impressive articles clearly written in a common mathematical language; they were available to readers of the XX century as well as readers of 1890-00s. We are trying to understand this phenomenon.
Sect. 5 contains three additions. Firstly, we briefly discuss historical sources on Alekseevsky and historiography.
Secondly, starting with [Ale1892], [Ale1894] Alekseevsky claimed many times over his researches could be extended to a wider class of functions depending on several parameters. We analyze these claims — clearly, he kept in mind the multiple gamma functions; definitely he did not publish theorems on this topic. However, these claims were known and made water run in circles. Apparently, [Bar1904] was a response to a declaration in [Ale1894].
To be complete, I present a list of his textbooks on calculus and mechanics published by universities (all research publications are listed in Bibliography, [Ale1884.1][Ale1904]; I think that this list is complete).
2. Biography of Alekseevsky
2.1. Early biography. Vladimir Petrovich Alekseevsky was born April 19 (Julian 07), 1858 in Novgorod governorate [^3] Russian: ‘губерния’ (‘guberniya’). and baptized in the village Peretënka located near Okulovka station on the railway line Moscow–St.Petersburg. His father, Petr Osipovich Alekseevsky, was a ‘praporshchik’ $\simeq$ warrant officer or sergeant major] of the Corps of transport engineers [Корпус инженеров путей сообщения], [Sin1917], [Gran2003].
In 1877, Alekseevsky graduated from the Nizhny Novgorod gymnasium. It was a good school, with many of its pupils going on to become well known, in particular, mathematicians Aleksandr Lyapunov and Ivan Privalov, physicist Grigory Landsberg [^4] Landsberg in particular could have won the Nobel in 1930 (combinational scattering of light), but was not nominated for it. He also edited the famous high-level ¡¡Elementary physics textbook¿¿, which was used by several generations of Russian natural scientists, engineers, and teachers., zoologist and ecologist Aleksandr Formozov, philosophers Semyon Frank and Vasily Rozanov, and chemist Alexey Favorsky.
2.2. Kharkov University, 1877-1905. After graduating from the gymnasium, Alekseevsky entered the Natural Science Department of the Faculty of Physics and Mathematics of Kharkov University. Two years later, he transferred to the Mathematical Department. In 1881, ‘for family reasons’, he left the university [^5] The main obstacle to his research activity was his difficult financial situation’ as was written in [Bakh1967]; this is also clear from Sintsov [Sin1917]). However, Sintsov in the obituary does not provide any details (and they are absent in other sources).. In 1884, he passed examinations without attending lectures and graduated from the university. At that moment he had three published papers [Ale1884.1]-[Ale1884.3] on the integrability of certain ordinary differential equations. They were considered as a graduate (‘candidate’) work [^6] There is no information about his scientific adviser(s). In [Ale1884.1]-[Ale1884.3], he mentioned the names of Vasily Imshenetsky (1832-1892), who worked at the Kharkov University 1877-1881 (and moved to St.-Petersburg afterward), and Aleksey Letnikov (1837—1888), who worked in Moscow..
Since 1881, Alekseevsky taught mathematics, mechanics, and physics in Kharkov in a real school [^7] A high school oriented to natural sciences, engineering, commerce, mathematics. (until 1891) and in the Second women’s gymnasium (until 1887). In 1887, he received a scholarship to prepare for a professorship (analogous to becoming a postgraduate student in the present).
@@ 2. Biography of Alekseevsky
Alekseevsky’s main work On functions similar to the gamma function was published in Communications of Kharkov Mathematical Society [^8] A well-known journal between 1879-1918. Russian: Сообщения и протоколы заседаний Математического общества при Императорском Харьковском университете; old Russian orthography: Сообщенiя и протоколы засданiй Математическаго общества при Императорскомъ Харьковскомъ университет; French: Communications et procus-verbaux de la société mathématique de Kharkow; German: Sammlung der Mitteilungen und Protokolle der Mathematischen Gesellschaft in Charkow. [Ale1888][Ale1889] (in fact, it is one paper split in the middle of a sentence). He defined the double gamma function $G(x;\alpha)$ satisfying two functional equations and derived an unexpectedly large collection of its properties. We discuss this impressive work and its genesis in Sect. 3.
In 1892, this work was presented as a dissertation [^9] A ‘magister dissertation’ in the terminology of that time. In modern terminology, it apparently corresponds to a PhD. [Ale1892]. After the defence (21.02.1893), he became a privatdozent in Kharkov University.
In March 1894 – November 1895 Alekseevsky visited Leipzig, Göttingen, and Paris. At that time, he published a short paper [Ale1894] in German in Leipz. Ber.. Apparently, his work became known due to this publication.
After coming back, he decided to leave the university and teached at a gymnasium in town Starobelsk, in Kharkov governorate since ’university …could do nothing to support the scientist it highly estimated’ (a quotation of a rapport of Vladimir Steklov in [Sin1917], who worked at Kharkov university in that time). Alekseevsky returned to the university the next year. Starting 1898, he also worked at Kharkov Technological Institute.
Four pages of his obituary [Sin1917] (and a big part of [StSi1904], see, also, the correspondence of Steklov and Aleksandr Lyapunov in [Ste1991]) contain a description of the efforts of several colleagues, through 1896-1904, to transfer Alekseevsky to the position of extraordinary professor. These included Aleksandr Lyapunov, Vladimir Steklov, Matvey Tikhomandritsky (specialist in elliptic functions), Dmitry Grave, Konstantin Andreev (later a professor and dean in Moscow), Ludwig Struve (astronomer [^10] Grandson of the famous astronomer Friedrich-Vasily Struve.), and Ivan Osipov (chemist). The nature of obstacles is not clear from the published literature. Steklov (letter to Lyapunov, 14.11.1903) characterized this as an ’absurd story’.
From [Sin1917]: Steklov [1902] explains, why V. P.’s doctoral dissertation $\simeq$ habilitation] was delayed – ’he could have compiled from the scientific results a fully satisfactory dissertation’ and expresses his conviction that ‘if V. P. did not take care to acquire formal rights, having at his disposal all the scientific materials, this circumstance can hardly diminish the scientific merits of V. P. Alekseevsky …But we must say even more. By now, V. P. Alekseevsky has prepared for printing and is already printing an extensive course of differential and integral calculus.…V. P. Alekseevsky’s course is not a template course adapted to the usual programs of higher educational institutions. It is an independent academic work, where, along with pedagogical requirements, it combines a rigorous and modern approach to the study of the subject’ (the book in question was not published).
After fifth attempt, 12.01.1904 Alekseevsky got the position of acting extraordinary professor in Kharkov. In the next year, he moved to Tomsk for a full professorship.
It seems doubtless (see Subsect. 3 and Subsect. 5) that Aleksevsky had a continuation of his work devoted to the multiple gamma functions and the multiple zeta functions. However, for unknown reasons he has not published these investigations. It seems also that this was one of the sources of his career difficulties (see footnote 23 below).
2.3. Tomsk Technological Institute, 1905–1916. The institute was founded in 1900 by chemist-technologist Efim Zubashev (see [Gal2004], [Gal2009]) and quickly became a center of education and science in Siberia. For instance, among its professors in the beginning of the XX century were Theodor Molien — one of the founders of representation theory of finite groups (together with George Frobenius and William Burnside) and theory of associative algebras [^11] See Bourbaki [Bou1960], Section ‘Noncommutative algebra’, van der Waerden [Wae1985], and selected works of Molien in [Mol1985]. — and the famous geologist Vladimir Obruchev.
According [Gran2003], Zubashev invited Alekseevsky to Tomsk for first time in 1900 (the Institute’s the first year of work). In April 1905, director Zubashev offered Alekseevsky the position of acting ordinary professor (full professor) of theoretical mechanics in Tomsk. Alekseevsky agreed and, starting 28.11.1905, worked in Tomsk. He left pure mathematics, and I could not find any his scientific publications by him of that time.
Granina’s paper [Gran2003] contains a lot of data on the educational and administrative activity of Alekseevsky in Tomsk. We mention only a few formal details.
Students of Tomsk Technological Institute were involved in the tumultuous events of the First Russian Revolution of 1905-07. Zubashev (who worked under pressure from both progressists and conservators) was accused of being loyal to the striking students and was exiled from Tomsk. Later, he returned to the director position but had to leave it ‘for health reasons’ (November 1907). The institute council elected Alekseevsky as the new director (rector), and his administrative activity was considered in local annals as quite successful and fruitful.
In 1911, there was a wave of student and professor protests in Russia against the education reform initiated by minister Lev Kasso. The Institute was involved in these protests. Several professors were dismissed for oppositionism [^12] The list of fired professors includes Molien and Obruchev. (8.06.1911). Aleksevsky and the institute council tried to reinstate them. As a result, by an imperial edict, Alekseevsky was ousted from the director position. However, he continued his full professorship and some administrative activity.
3. Alekseevsky’s work, 1888-89
Alekseevsky’s work On functions similar to the gamma function was published in Communications of Kharkov Mathematical Society in 1888-1889 in two papers [Ale1888], [Ale1889].
@@ 3. Alekseevsky’s work, 1888-89
3.1. The function $G(x)$. The Alekseevsky’s paper begins with the words: The first problem, whose solution is the topic of the present investigation, is to examine the properties of a function $G(x)$ satisfying the equation:
@@
There are no references, no names, no motivations, no acknowledgements here. It is hard to believe that the author simply decided to attack such an exotic problem and successfully solved it. In reality, the presentation of the paper was backwards to the real way of thinking; this is discussed below in Subsect. 3.
Taking the logarithm of both sides of (3.1), we come to the difference equation
@@
According to the Malmsten formula, see for example [Erd1953], formula 1.9(1),
@@
Denote the integrand by $L_{1}(x,u)$. Alekseevsky (§1) writes the simplest function $L_{2}(x,u)$ such that
@@
is the Euler constant (cf. the similar expression for $\Gamma(x)$ in [Erd1953], 1.1 (3)). Next, he notices that this formula gives an expression for $G(x)$ on the whole complex plane.
@@
After complicated calculations with functions $G(x;\alpha)$ in §21 (see below Subsect. 3) he gets the multiplication formula (cf. [Erd1953], 1.2(11)) in the form
@@
The multiplication formula admits a simpler form obtained by Kinkelin [Kin1860], see formula (4.2) below.
In §11, Alekseevsky establishes the following counterpart of the ’Stirling’ asymptotic expansion (cf. [Erd1953], 1.18(1)):
@@
In §16, he introduces functions $G_{n}(x)$ satisfying the relations
@@
which implies the desired property.
Alekseevsky does not develop a further theory of such functions, but in §17-19 he discusses some generalities of function theory related to the Weierstrass factorization theorem. Namely, he considers arbitrary products of the form
@@
He represents $\ln f_{0}(z)$ as a Laplace transform, and, applying the same trick with functions $P_{n}(z)$, constructs a sequence of functions $f_{1}(z)$, $f_{2}(z)$, …such that
$f_{n}(z+1)=f_{n}(z)f_{n+1}(z).$
He writes explicit representations of functions $f_{n}$ as products of functions $G_{m}$ with given $m\leqslant n$.
3.4. Double gamma functions with two periods. In §20, Alekseevsky writes: Functions similar to $\Gamma(x)$ considered above are special cases of more general functions. Since they are related to theta-functions, we derive some properties of one such function.
Fix a parameter $\alpha$ (which initially is positive real, but later it becomes complex). Alekseevsky considers a function [^13] He uses the notation $H(x)$ or $H(x,\alpha)$. $G(x;\alpha)$ such that
@@
Alekseevsky finds the main properties (cf. [AlKu2023]) of this function, in particular,
@@
see §21. This formula has numerous interesting special cases, for instance, we can set $m=1$, or $n=1$, or $\alpha=1$, or $\alpha=1/n$, or $\alpha=n$, or $m=n$. The multiplication formula (3.8) also is a non-obvious corollary of the last formula.
@@
It is clear that this function consists of a quarter of the factors entering to the $\varsigma_{3}$-function of Weierstrass [^14] The Weierstarss function, see for example [Akh1970], §13, is given by a product (3.19) over $(m,n)\in\mathbb{Z}^{2}\setminus(0,0)$. In our case (3.16), the product is given over $m\geqslant 0$, $n\geqslant 0$, $(m,n)\neq(0,0$. (or $\vartheta_{1}$ of Jacobi), herein it comes from $\varsigma_{3}$ the same way as $\frac{1}{\Gamma(z)}$ comes from $\sin\pi z$. However, in a two-periodic function, the ratio $\frac{\omega^{\prime}}{\omega}$ can be imaginary or complex but for the function $G$ this condition is not necessary151515i.e., a ratio of periods can be positive real..
The function $G$ is analogous to the function of Heine [^16] Footnote from [Ale1888]: Handbuch der Kugelfunctionen, p. 109, or Crelle J. t 34. p. 290, see [Hei1847], [Hei1878]. I do not understand these references (they are taken from [App1881]). However, the attribution of this expression to Eduard Heine is correct, see the next subsection., which is composed from a half of the factors of $\vartheta_{3}$. For a collation, we present an expression from Appell’s memoire [^17] Footnote from Alekseevsky: Mathematische Annalen. B. 19. p. 84. , see [App1881],
@@
This explains the origin of this work. Looking to the formula in Appell’s paper, Alekseevsky understood that it is possible to manipulate with Weierstrass-type product over a quadrant in the lattice $\mathbb{Z}^{2}$. However, he found that the integral representation (3.4) is the key to the whole story, and the initial definition of the work became the final theorem of his published paper (in fact, an evaluation of constants $a$, $b$ in (3.16) is a highly nontrivial problem).
The work discussed here became Alekseevsky’s dissertation [Ale1892]. It was published as a separate book and is a literal reprint of [Ale1888]-[Ale1889] with of 4 pages added to the beginning. The introduction contains the following sentence:
The known dependence between the Euler function of the second kind [the gamma function] and the sine [i.e., the reflection formula $\Gamma(z)\Gamma(1-z)=\pi/\sin(\pi z)$] made me think to extract from the function $\vartheta_{1}$ of Jacobi a function similar to $1/\Gamma(x)$ and investigate its properties.
By way of analogy with the function $\Gamma(x)$, I came to consider functions whose roots have the form $m\omega+n\omega^{\prime}$, where $m$ and $n$ are negative integers, and $\omega$, $\omega^{\prime}$ are periods of elliptic functions.
As we have seen above, Alekseevsky introduced additional parameters $\omega_{1}$, $\omega_{2}$ to the function $G=G_{2}(x)$. After Barnes [Bar1904], we know that it is possible to introduce similar parameters $\omega_{1}$, …, $\omega_{n}$ to the functions $G_{n}(x)$ and get ‘multigamma functions’. Such functions (we use a non-standard notation) satisfy the collection of identities
@@
It is strange that Appell himself did not introduce the double gamma. In any case, Alekseevsky’s response to Appell’s paper was completely natural.
@@
2. Roots of functions similar to $\Gamma$ have the form